In an office building, a room contains 90 chairs. The number of chairs per row is three less than three times the number of rows. Find the number of rows and the number of chairs per row.

Found 2 solutions by edjones, ankor@dixie-net.com:
Answer by edjones(7569)

Let r=number of rows, c=number of chairs/row
cr=90
c+3=3r
.
c=3r-3
r(3r-3)=90
3r^2-3r-90=0
r^2-r-30=0
(r-6)(r+5)=0
r=6
c=15
.
Ed

Answer by ankor@dixie-net.com(15649)

In an office building, a room contains 90 chairs.
The number of chairs per row is three less than three times the number of rows.
Find the number of rows and the number of chairs per row.
:
Let r = no. of rows
Let n = no. of chair/row
:
r * n = 90
:
"The number of chairs per row is three less than three times the number of rows."
n = 3r - 3
:
Replace n in the equation: r*n = 90
r*(3r-3) = 90
3r^2 - 3r = 90
3r^2 - 3r - 90 = 0
simplify, divide by 3
r^2 - r - 30
Factor
(r-6)(r+5) = 0
positive solution
r = 6 rows
then
n = 3(6) - 3
n = 15 chairs in each row
:
:
Check: 15 * 6 = 90